OpenStudy (anonymous):

explain in complete sentences how you would use the substitution method to solve the following system of equations and provide the solution to the system . x+5y=3 2x+9y=4

OpenStudy (anonymous):

put x=3-5y in second equation and solve it

OpenStudy (anonymous):

2(x+5y-3) 2x+10y-6 -- --- -- 2 2 2 x = 5y -3

OpenStudy (anonymous):

no first you have too multiple the first equation by (2) and the second by (-1) so now you have a positive 2x and a -2x it will look like this 2x+10y=6 =-2x-9y=-4 the 2x cancels out which leaves 1y/1=2/1 y=2 then you just choose which equation you want too substitute y for and solve

OpenStudy (anonymous):

and that equal sign was suppose too be a (+) sign sorry